# Holographic homogeneous superfluid on the sphere

### v1

@@ Abstract
In this paper, we extend the study of holographic superfluids from planar topology to spherical topology, inspired by recent studies on Bose-Einstein condensation (BEC) superfluid on shell-shaped geometry. We mainly focus on superfluid phase transition and Quasi Normal Modes (QNMs). It turns out that the critical temperature for the superfluid phase transition on the sphere is higher than that in the planar case. We investigated four different solutions in the backgrounds of large black hole and small black hole. The calculation of free energy selects the most stable solution. Finally, after calculating the quasi-normal modes and their dynamic behavior, we obtained three different channels similar to the planar superfluid case, along with the ”first” hydrodynamic excitation, which maybe is associated with Berezinskii-Kosterlitz-Thoules (BKT) transition.
@@ 1 Introduction and motivation
The exploration of superfluidityONSAGER ; Transport phenomena , a state of matter characterized by zero viscosity and perfect fluidity without dissipating energy, has been a fascinating subject in the realm of condensed matter physics. Traditionally, superfluids have been studied in the context of Bose-Einstein condensates (BECs), where the phenomenon is observed at extremely low temperaturesLondon ; Anderson ; Davis ; GP ; BEC-Superfluidity ; Nonlinear Phenomena . Since then, the theoretical framework and experimental techniques for studying superfluids have been thriving, leading to novel insights and applications in various fields, including quantum information processingQuantum copying ; Bose metal ; Quantum Simulation , superconductivitySuperconductivity ; Superconductivity1 , and astrophysicsNeutron Star ; Neutron Star1 . In recent years, there has been a surge of interest in studying superfluids on shell-shaped systemsBEC-sphere ; Shell-shaped Atomic ; Bereta , inspired by the unique properties that arise from the curvature and topology of the confining geometryObservation ; microgravity ; microgravity1 ; orbit . This shift is motivated by the need to understand superfluid behavior in environments that more closely resemble real-world physical systems, which often do not conform to simple planar geometries. Shell-shaped systems offer a rich playground for investigating the interplay between superfluid properties and the geometry structure, potentially leading to new picture of phase transition and collective excitations Hollow ; Static and dynamic ; Topological superfluid that are not present in the planar case.
@@
The holographic principle, or AdS/CFT correspondence, has emerged as a powerful tool in theoretical physics, allowing the study of strongly coupled systems through the gravity duals in higher-dimensional spacetimeSusskind ; Maldacena ; Witten . The study of superfluidity in holographic systems has garnered significant interest due to its implications for understanding strongly correlated systems and quantum phase transitions, such as holographic superfluid key-9 ; key-10 ; key-11 ; key-12 ; Guo ; gaomeng and superconductorsNishioka , where the phase transitions and properties of these systems can be mapped to black hole solutions in Anti-de Sitter (AdS) spacetime. Despite considerable advances in the comprehension of holographic superfluids within planar topology, the extension to spherical topology remains notably unexplored. This paper seeks to address this gap by exploring the phase transitions and collective modes of holographic superfluids within a shell-shaped conformal boundary. Our motivation stems from the desire to understand how the spherical geometry affects the superfluid phase transition, the stability of the superfluid phase, and the dynamics of collective excitation modes.
@@
The discussion in reference Kaku includes holographic superconducting phase and linear perturbations analysis under spherical topology, however, it primarily addresses the scenario of small black holes, which are inherently thermodynamically unstable spacetime structures. Their study does not explore the critical phase transition temperature affected by the geometric topological structure. Furthermore, they just focus on the case of high frequency $\omega$, ignoring other intriguing collective modes for small frequency case. Our research is based on the Abelian-Higgs model in a Schwarzschild-AdS4 spacetimeS.A ; C.P , which provides a simple yet rich framework to analyze superfluid properties. From our exploration, we find it has a larger critical temperature for the phase transition from normal fluid to the superfluid on the unit sphere than that on the plane. In our study, there exist four different solutions: hairy large black hole, bold large black hole, hairy small black hole, and bold small black hole. For the four distinct solutions, we analyzed their phase-transition behavior and calculated free energy. On the other hand, we calculated the quasi normal modes as well as its dynamic behavior associated with temperature, chemical potential and angular quantum number. Moreover we obtain the ”first” hydrodynamic excitation mode and three different channels of modes, similar to those on the plane topology.
@@
The paper is organized as follows. In the next section, we introduce the holographic superfluid setup under a spherical topology background. In Section III, we study the phase transition from normal state to superfluid state. In Section IV, we perform a linear analysis to explore collective excitation modes. In Section V, we draw our conclusions and some outlook.
@@ 2 Holographic setup
In the following, we shall work in the units with $L=1$, then the above temperature will be interpreted as the temperature of the dual boundary system living on the unit sphere by holography. Obviously, at $z_{h}=\sqrt{3}$, Hawking temperature arrives at its minimal value $T_{min}=\frac{\sqrt{3}}{2\pi}$, below which there is no black hole solution and above which there are two black hole solutions. Moreover, there exists a first order Hawking-Page phase transition at $T_{HP}=\frac{1}{\pi}$, above which the thermodynamically stable state is given by the large black hole and below which the thermodynamically stable state is given by the thermal AdS HP . However, the large black hole is dynamically stable at least at the linear level all the way down to the minimal temperature $T_{min}$. Similarly, although the small black hole is thermodynamically unstable, it is also dynamically stable at least at the linear level all the way down to $T_{min}$. So in what follows, we would like to take both large and small black holes as our background geometry, regardless of whether the background geometry is thermodynamically stable or not.
@@
with the renormalized action $S_{ren}=S-\int d^{3}x\sqrt{-h}|\Psi|^{2}$uo . When the scalar source is turned off, $\langle O\rangle\neq 0$, the bulk black hole carries a scalar hair, corresponding to the superfluid state on the boundary with $\langle O\rangle$ interpreted as the superfluid condensate, otherwise it is bold, corresponding to the normal fluid state.
@@ 3 Holographic superfluid phase transition
In this section, we shall investigate the phase transition to holographic static and homogeneous superdluid from the normal fluid. As such, the non-vanishing bulk fields can be assumed to be $\Psi(z)\equiv z\psi(z)$ and $A_{t}(z)$ with $\psi(z)$ being also real. Then the corresponding equations of motion reduces to
@@
Next we present the typical condensation behaviors across the phase transition point in Figure 3 and Figure 4 at fixed temperature and chemical potential, respectively. As shown in Figure 3, the condensation displays a normal second order phase transition with respect to the chemical potential at fixed temperature. In particular, this behavior is universal in the sense that neither does it depend on whether the temperature under consideration is higher or lower than the Hawking-Page temperature, nor depends on whether the involved background geometry is given by the large or small black hole. On the other hand, as demonstrated in Figure 4, although the condensation associated with the large black hole as the background geometry still exhibits a normal second phase transition with respect to the temperature at fixed chemical potential, the condensation associated with the small black hole as the background geometry demonstrates an anomalous second order phase transition in the sense that below some critical temperature, the condensate instead vanishes while above the critical temperature, the condensate increases with the increase of the temperature.
@@
With it, we plot in Figure 5 the resulting difference of the free energy of all other possible phases with respect to that for the bold small black hole at given temperature and chemical potential. As one can see from the free energy difference, the phase transition indicated in Figure 3 and Figure 4 is actually a second phase transition. In addition, the hairy large black hole, if such a solution exists, always carries the lowest free energy, thus thermodynamically favorable. Otherwise the lowest free energy carrier is the bold large black hole.
Figure 5: The difference in the free energy between all other possible phases and the bold small black hole. The dashed horizontal line denotes the free energy for the bold small black hole, which we set to be zero. the blue, dashed red, and red lines individually denote the free energy of hairy small black hole, bold large black hole, and hairy large black hole with respect to that for the bold small black hole.
@@ 4 Collective modes of holographic superfluid
It is noteworthy that Eq. (25) and Eq. (26) should be discarded in the case of $l=0$, where there is essentially no excitation for the $b$ and $c$ modes indeed, which can be seen from Eq. (15) and Eq. (16). Eqs. (22-26) in the bulk as well as Eq. (27) evaluated at the AdS boundary supplemented with the Dirichlet boundary conditions for all the modes on the AdS boundary can be formulated in terms of the generalized eigenvalue problem for the quasi-normal modes, which can be solved readily by our numerics. Furthermore, note that the $b$ mode decouples from the other modes. So in the case of $l=0$, we have the quasi-normal frequencies from the coupled $q_{1}$, $q_{2}$, $a$ modes, while in the case of $l\neq 0$ we have the quasi-normal frequencies separately from the $b$ mode and the coupled $q_{1}$, $q_{2}$, $a$, $c$ modes, respectively.
@@
We first look at the behavior of the $l=0$ quasi-normal modes near the phase transition, which is depicted in Figure 6 and Figure 7. As the normal fluid approaches the critical point, the two low-lying modes, which lie symmetrically in the lower $\omega$ plane, migrate symmetrically towards the origin. Exactly at the critical point, they meet each other at the origin. If the system is kept in the normal fluid state beyond the critical point, one can see the two low-lying modes will keep migrate towards the upper $\omega$ plane, which signals the dynamical instability of the normal fluid state. On the other hand, if the system transits to the superfluid state, one can see that one mode remains at the origin as the Goldstone mode while the other mode migrates down the imaginary axis as the Higgs mode. Although our system is confined on the finite sphere, the above result indicates the equivalence of the dynamical instability and thermodynamical instability still holds for our system.
@@
Figure 6: The variation of the low-lying quasi-normal modes with the chemical potential at fixed temperature. The dashed red and dashed blue lines are for the bold large and small black holes while the red and blue lines are for the hairy large and small black holes. The chemical potential is varied from 3.0 to 5.5 for the top left panel, from 1.8 to 4.4 for the top right panel, from 5.4 to 7.8 for the bottom left panel, and from 1.4 to 4.0 for the bottom right panel.
@@
Figure 7: The variation of the low-lying quasi-normal modes with the temperature at fixed chemical potential. The dashed red and dashed blue lines are for the bold large and small black holes while the red and blue lines are for the hairy large and small black holes. The temperature is varied from 0.3 to 0.5 for the left panel, and from 0.29 to 0.41 for the right panel.
@@
In what follows, we investigate the behavior of the superfluid system for the case of $l\neq 0$ with large black hole and small black hole as our background geometry. Initially, we studied the variation of the low-lying quasi-normal modes with the temperature at fixed chemical potential, which is presented in Figure 8. The two top panels are calculated in the background of large black hole, the bottom two panels are from small black hole. From the top left panel, two scalar modes move towards the origin with temperature decreasing, however, they change direction after finishing the phase transition represented by the green dot. Until the temperature reaches the lowest temperature (characterized by purple dot), both modes remain stable. However, the scalar modes in the small black hole (bottom left panel) exhibit anomalies with changes in temperature. When the temperature is low, the system is in a normal flow state, while at higher temperatures, the system transits to a superfluid state. The panels of right column is from the independent equation of the $b$ field. It exists an inflection point at the location of the phase transition for large black hole, but no inflection point appears in the case of a small black hole. Meanwhile, we explore the impact associated with chemical potential on the low-lying quasi-normal modes at fixed temperature, which is shown in Figure 9. Similarly, the dynamic behavior of the low-lying quasi-normal modes undergoes a turning point at the phase transition. Remarkably, the low-lying quasi-normal modes of $b$ field do not move until the phase transition occurs, which is reasonable according to Equation (25).
@@
Figure 8: The variation of the low-lying quasi-normal modes with the temperature at fixed chemical potential for large black hole (top plots) and small black hole (bottom plots), where angular quantum number is set to $l=1$. The temperature is varied from $T_{min}=\sqrt{3}/2\pi$ to 0.52 for large black hole, from $T_{min}=\sqrt{3}/2\pi$ to 0.743 for small black hole. As the temperature decreases, the two modes move in the direction indicated by the arrow. The left plots are the case of scalar channel and the right plots are from independent equation of $b$ field. The green dot represents phase transition from normal state to superfluid state for large black hole, the situation is exactly opposite for small black hole. The purple dot represents the state of the lowest temperature.
@@
Figure 9: The variation of the low-lying quasi-normal modes with chemical potential at fixed temperature for $l=1$ case under large black hole (top plots) and small black hole (bottom plots) background. Chemical potential is varied from 3.5 to 5.0 for large black hole, from 2.2 to 3.5 for small black hole. Left plots are the case of scalar channel and the right plots are from independent equation of $b$ field. The green dot represents phase transition from normal state to superfluid state for both two types of black holes. Thus, we can identify that the modes form $b$ field do not move until phase transition occurs.
@@
Furthermore, we investigate the impact of angular quantum numbers $l$ on the collective modes. Figure 10 shows the dynamic behavior of the collective modes with respect to the angular quantum number at fixed temperature and chemical potential within normal-fluid state. Actually, there exist three different channels in homogeneous superfluid: the scalar field channel, the transverse gauge field channel and the longitudinal gauge field channel three channels . The three kinds of channels are decoupled when the system is in normal-fluid state. In Figure 10, the three top panels are obtained on large black hole background, bottom results are from small black hole. The left column of panels is about scalar field channel ($q_{1}$ and $q_{2}$ fields), middle column of panels is about longitudinal gauge field channel (coupled $a$ and $c$ fields), the right column of panels is from transverse gauge field ($b$ field) channel. $l$ is varied from 0 to 6 for the scalar channel and from 1 to 6 for gauge field channels. Since the fields of $b$ and $c$ are excitation modes of the angular components of the gauge fields, these excitation modes do not exist as $l=0$. Therefore, we calculate QNMs of $b$ field and $c$ field starting from $l=1$. Moreover, $a,c$ fields are coupled as $l\neq 0$ for normal-fluid state. There is only one zero mode for $a$ field, which reflects charge diffusion. Since $L=1$ on the sphere topology corresponds to the case of large wave number $k$ on the planar topology, the dynamic behaviors of the three channels for large black hole in Figure 10 are similar to those in the planar case. However, we can identify that the dynamic behaviors of scalar modes in the backgrounds of the two types of black holes are completely different. Here, small black holes also exhibits an unusual side. Therefore, compared to small black hole, the superfluid system with large black hole being background geometry in spherical topology is similar to the case in planar topology. Figure 11 describes the superfluid case for different values of $l$. The left column is from scalar modes, right column is about transverse gauge field. In this case, scalar field channel and longitudinal gauge field channel are coupled together, the transverse gauge field channel is always decoupled with other fields. Similarly, the dynamic behaviors of modes under large black hole background in Figure 11 are consistent with those of the planar case, but there are anomalies in the case of small black holes.
@@
Figure 10: The variation of the low-lying quasi-normal modes with angular quantum number $l$ at fixed temperature and chemical potential within normal-fluid state for large black hole (top panels) and small black hole (bottom panels). The leftmost panels are about scalar field channel ($q_{1}$ field and $q_{2}$ field), middle panels are about longitudinal gauge field channel ($a$ field and $c$ field), the rightmost panels are about transverse gauge field ($b$ field). $l$ is varied from 0 to 6 for leftmost panels, from 1 to 6 for other panels.
@@
Figure 11: The variation of the low-lying quasi-normal modes with angular quantum number $l$ at fixed temperature and chemical potential within superfluid state for large black hole (top plots) and small black hole (bottom plots). The left plots are from scalar field modes, right plots are from transverse gauge field. $l$ is varied from 0 to 6 for top left panel, from 1 to 6 for other three panels.

### v4

@@ Abstract
In this paper, we extend the study of holographic superfluids from planar topology to spherical topology, inspired by recent studies on Bose-Einstein condensation (BEC) on shell-shaped geometry. We investigate the superfluid phase transition from normal fluid and its Quasi-Normal Modes (QNMs) on the sphere. It turns out that the critical temperature for the superfluid phase transition on the sphere is higher than that in the planar case. We investigated four different solutions in the backgrounds of large and small black holes. The calculation of free energy selects the most stable solution. Finally, after calculating the quasi-normal modes and their dynamic behavior, we obtained three different channels similar to the planar superfluid case, along with the “first” hydrodynamic excitation mode.
@@ 1 Introduction and motivation
The exploration of superfluidity ONSAGER ; Transportphenomena , a state of matter characterized by zero viscosity and perfect fluidity without dissipating energy, has been a fascinating subject in the realm of condensed matter physics. Traditionally, superfluids have been studied in the context of Bose-Einstein condensates (BECs), where the phenomenon is observed at extremely low temperatures London ; Anderson ; Davis ; GP ; BEC-Superfluidity ; NonlinearPhenomena . Since then, the theoretical framework and experimental techniques for studying superfluids have been thriving, leading to novel insights and applications in various fields, including quantum information processing Quantumcopying ; Bosemetal ; QuantumSimulation , superconductivity Superconductivity ; Superconductivity1 , and astrophysics NeutronStar ; NeutronStar1 . In recent years, there has been a surge of interest in studying superfluids on shell-shaped systems Shell-shapedAtomic ; Bereta ; BEC-sphere , inspired by the unique properties that arise from the curvature and topology of confining geometry Observation ; microgravity ; microgravity1 ; orbit . In the field of condensed matter, Ref.Bereta theoretically studied the critical temperature of Bose-Einstein condensation in a spherical potential well and discussed the effect of dimensional changes of the spherical potential well, from 3D to 2D, on the critical temperature. However, they cannot rigorously obtain a finite critical temperature for Bose Einstein condensation on the 2D sphere using the semi-classical approximation method. Moreover, as the external radius $b=a+\delta$ goes to the internal radius $a$, $\delta\rightarrow 0$, Taylor expansion approximation about Bessel function gives the conclusion that it need an infinite amount of energy to excite the radial degree of freedom. Therefore, the so-called 2-dimensional sphere they refer to is not a strictly 2-dimensional spherical system. In contrast, the holographic gravity model S.A ; C.P in this paper can easily and analytically construct a mathematical 2-dimensional spherical system to study the phase transition of Bose-Einstein condensation, where the thickness of its shell is strictly zero. Ref.BEC-sphere investigated Bose-Einstein condensation in spherical systems, which includes two cases: noninteracting Bose gas and interacting Bose gas. In the non-interacting Bose gas system, they found that the smaller the radius of the spherical surface, the higher the critical temperature for Bose-Einstein condensation. Furthermore, the critical temperature for the phase transition on the spherical surface is always higher than that in the planar system, which is mirrored in our paper. In addition to theoretical research, there have also been some experimental advancements. A spherical potential well can be constructed experimentally PhysRevLett.86.1195 ; PhysRevA.69.023605 ; Garraway_2016 to confine the atoms in this kind of bubble trap. Recently, with the construction of the NASA Cold Atom Lab (CAL) inside the International Space Station Aveline2020 , experiments involving the cold atoms system in a bubble trap under microgravity conditions have become possible Lundblad_2019 . Furthermore, ultracold bubbles between two-component cold atomic gases has been observed in microgravity aboard CAL Observation , which confines the Bose-Einstein condensation formed by one component, restricting its exclusion to the surface shell of the solid sphere of the Bose-Einstein condensate formed by the other component. Ultracold bubbles are precisely spherical BEC systems, and the spherical holographic superfluid model has the potential to describe ultracold bubbles. The two-dimensional surface will exhibit richer topological structures and curvature behaviors and offer a rich playground for investigating the interplay between superfluid properties and the geometry structure, potentially leading to a new picture of phase transition and collective excitations Hollow ; Staticanddynamic ; Topologicalsuperfluid that are not present in the planar case. Therefore, the study of spherical superfluids is of great significance, and the results are expected to be experimentally validated.
Inspired by the progress in these experiments and the theoretical research about shell-shaped Bose Einstein condensations in the realm of condensed matter, we have begun to use the holographic gravity approach to study the Bose-Einstein condensation of cold atoms in a two-dimensional spherical system to explore the new phenomena that may be induced by the spatial topology and curvature within it. One of our results shows that the superfluid phase transition temperature on the 2-dimensional unit sphere is indeed higher than that in the planar topology case, which is consistent with the result in Ref.BEC-sphere . As for the case of interacting Bose gases in Ref.BEC-sphere , it can be realized by adding interaction terms for the scalar field in a holographic gravity model, but this is not the focus of this article.
@@
The holographic principle, or AdS/CFT correspondence, has emerged as a powerful tool in theoretical physics, allowing the study of strongly coupled systems through gravity duals in higher-dimensional spacetime Susskind ; Maldacena ; Witten . The study of superfluidity in holographic systems has garnered significant interest due to its implications for understanding strongly correlated systems and quantum phase transitions, such as holographic superfluid key-9 ; key-10 ; key-11 ; key-12 ; Guo ; gaomeng and superconductors Nishioka , where the phase transitions and properties of these systems can be mapped to black hole solutions in Anti-de Sitter (AdS) spacetime. Despite considerable advances in the comprehension of holographic superfluids within planar topology, the extension to spherical topology remains notably unexplored. This paper seeks to address this gap by exploring the phase transitions and collective modes of holographic superfluids on the sphere, and plans to understand how the spherical geometry affects the superfluid phase transition as well as the dynamics of collective excitation modes.
@@
Furthermore, The discussion in Kaku includes holographic superconducting phase and linear perturbation analysis under spherical topology; however, it primarily addresses the scenario of small black holes, which are inherently thermodynamically unstable spacetime structures. Moreover, they just focus on the case of high frequency $\omega$, ignoring other intriguing collective modes for the case of low frequency. In our study, we both focus on the background geometry from large black hole and small black hole, and investigate collective elementary excitation coming from small $\omega$ and compare them with the flat case. There exist four different solutions: hairy large black hole, bald large black hole, hairy small black hole, and bald small black hole. For the four distinct solutions, we analyzed their phase-transition behavior and calculated their free energy. On the other hand, we calculated the quasi-normal modes as well as its dynamic behavior associated with temperature, chemical potential, and angular quantum number three parameters. Moreover, we obtain the “first” hydrodynamic excitation mode, which is reflected in the literature Topologicalsuperfluid . Meanwhile, we observed three kinds of quasi-normal modes coming from different channels.
@@
The paper is organized as follows. In the next section, we introduce the holographic superfluid setup under a spherical topology background. In Sec.III, we study the phase transition from normal state to superfluid state. In Sec.IV, we perform a linear analysis to explore collective excitation modes. In Sec.V, we draw our conclusions and some outlook.
@@ 2 Holographic setup
In the following, we shall work in the units with $L=1$, then the above temperature will be interpreted as the temperature of the dual boundary system living on the unit sphere by holography. Note that, As usual, in this paper, we adopt natural units ($c=1,G=1,\hbar=1$), where all physical quantities have the dimension of length. Additionally, for convenience, we set the AdS radius to 1 ($L=1$), so all physical quantities are dimensionless. Obviously, in natural units, at $z_{h}=\sqrt{3}$, Hawking temperature arrives at its minimal value $T_{min}=\frac{\sqrt{3}}{2\pi}$, below which there is no black hole solution and above which there are two black hole solutions. Moreover, there exists a first order Hawking-Page phase transition at $T_{HP}=\frac{1}{\pi}$, above which the thermodynamically stable state is given by the large black hole and below which the thermodynamically stable state is given by the thermal AdS HP . However, the large black hole is dynamically stable at least at the linear level all the way down to the minimal temperature $T_{min}$. Similarly, although the small black hole is thermodynamically unstable, it is also dynamically stable at least at the linear level all the way down to $T_{min}$. So in what follows, we would like to take both large and small black holes as our background geometry, regardless of whether the background geometry is thermodynamically stable or not.
@@
with the renormalized action $S_{ren}=S-\int d^{3}x\sqrt{-h}|\Psi|^{2}$ Guo . When the scalar source is turned off, $\langle O\rangle\neq 0$, the bulk black hole carries a scalar hair, corresponding to the superfluid state on the boundary with $\langle O\rangle$ interpreted as the superfluid condensate, otherwise it is bald, corresponding to the normal fluid state.
@@ 3 Holographic superfluid phase transition
In this section, we shall investigate the phase transition to holographic static and homogeneous superfluid from the normal fluid. As such, the non-vanishing bulk fields can be assumed to be $\Psi(z)\equiv z\psi(z)$ and $A_{t}(z)$ with $\psi(z)$ being also real. Then the corresponding equations of motion reduces to
@@
Next we present the typical condensation behaviors across the phase transition point in Figure 3 and Figure 4 at fixed temperature and chemical potential, respectively. As shown in Figure 3, the condensation displays a normal second order phase transition with respect to the chemical potential at fixed temperature. In particular, this behavior is universal in the sense that neither does it depend on whether the temperature under consideration is higher or lower than the Hawking-Page temperature, nor depends on whether the involved background geometry is given by the large or small black hole. On the other hand, as demonstrated in Figure 4, although the condensation associated with the large black hole as the background geometry still exhibits a normal second phase transition with respect to the temperature at fixed chemical potential, the condensation associated with the small black hole as the background geometry demonstrates an anomalous second-order phase transition in the sense that below some critical temperature, the condensate instead vanishes, while above the critical temperature, the condensate increases with the increase of the temperature.
@@
With it, we plot in Figure 5 the resulting difference of the free energy of all other possible phases with respect to that for the bald small black hole at given temperature and chemical potential. As is well known, the equilibrium at fixed temperature is unstable for the lower mass of black hole but is locally stable for the higher mass of black hole HP . As such the free energy of large black holes is lower than that of small black holes. In this paper, we only calculated the free energy of the matter field in two different black hole backgrounds. The study finds that the free energy of the matter field in the large black hole background is smaller than that in the small black hole background. In conclusion, the total free energy of matter and gravitational fields in the large black hole spacetime background is lower than that in the small black hole case, which indicates that the physical system in the large black hole background is more thermodynamically stable. On the other hand, in this article, we focus on the probe limit case, therefore, if we consider the contribution of gravitational free energy and matter together, the free energy of the matter field can be neglected, which has no significance for the present work. Here, we use the small bald black hole as a reference, because we can directly compare the thermodynamic stability of small and large black holes, and we are not concerned with their absolute values, only with their relative sizes, so we can determine which one is more stable. The results indicates that we will focus only on the case of large black holes in our future studies.
As one can see from the free energy difference, the phase transition indicated in Figure 3 and Figure 4 is actually a second-order phase transition. In addition, the hairy large black hole, if such a solution exists, always carries the lowest free energy, thus thermodynamically favorable. Otherwise, the lowest free energy carrier is the bald large black hole.
Figure 5: The difference in the free energy between all other possible phases and the bald small black hole. The dashed horizontal line denotes the free energy for the bald small black hole, which we set to be zero. The blue, dashed red, and red lines individually denote the free energy of hairy small black hole, bald large black hole, and hairy large black hole with respect to that for the bald small black hole.
@@ 4 Collective modes of holographic superfluid
It is noteworthy that Eq. (25) and Eq. (26) should be discarded in the case of $l=0$, where there is indeed essentially no excitation for the $b$ and $c$ modes, which can be seen from Eq. (15) and Eq. (16). Eqs. (22-26) in the bulk as well as Eq. (27) evaluated at the AdS boundary supplemented with the Dirichlet boundary conditions for all modes on the AdS boundary can be formulated in terms of the generalized eigenvalue problem for the quasi-normal modes, which can be solved readily by our numerical method. Furthermore, note that the $b$ mode decouples from the other modes. So in the case of $l=0$, we have the quasi-normal frequencies from the coupled $q_{1}$, $q_{2}$, $a$ modes, while in the case of $l\neq 0$ we have the quasi-normal frequencies separately from the $b$ mode and the coupled $q_{1}$, $q_{2}$, $a$, $c$ modes, respectively.
@@
We first look at the behavior of the $l=0$ quasi-normal modes near the phase transition, which is depicted in Figure 6 and Figure 7. As the normal fluid approaches the critical point, the two low-lying modes, which lie symmetrically in the lower $\omega$ plane, migrate symmetrically towards the origin. Exactly at the critical point, they meet each other at the origin. If the system is kept in the normal fluid state beyond the critical point, one can see that the two low-lying modes will continue to migrate towards the upper $\omega$ plane, which signals the dynamical instability of the normal fluid state. On the other hand, if the system transits to the superfluid state, one can see that one mode remains at the origin as the Goldstone mode, while the other mode migrates down the imaginary axis as the Higgs mode. Although our system is confined on the finite sphere, the above result indicates the equivalence of the dynamical instability and the thermodynamical instability still holds for our system.
@@
Figure 6: The variation of the low-lying quasi-normal modes with the chemical potential at fixed temperature. The dashed red and dashed blue lines are for the bald large and small black holes while the red and blue lines are for the hairy large and small black holes. The chemical potential is varied from 3.0 to 5.5 for the top left panel, from 1.8 to 4.4 for the top right panel, from 5.4 to 7.8 for the bottom left panel, and from 1.4 to 4.0 for the bottom right panel.
@@
Figure 7: The variation of the low-lying quasi-normal modes with the temperature at fixed chemical potential. The dashed red and dashed blue lines are for the bald large and small black holes while the red and blue lines are for the hairy large and small black holes. The temperature is varied from 0.3 to 0.5 for the left panel, and from 0.29 to 0.41 for the right panel.
@@
In what follows, we investigate the behavior of the superfluid system for the case of $l\neq 0$ with large black hole and small black hole being our background geometry, respectively. Initially, we studied the variation of the low-lying quasi-normal modes with the temperature at fixed chemical potential, which is presented in Figure 8. The two top panels are calculated in the background of the large black hole, the bottom two panels are from small black hole. From the top left panel, two scalar modes move towards the origin with temperature decreasing; however, they change direction after finishing the phase transition represented by the green dot. Until the temperature reaches the lowest value (characterized by purple dot), both modes remain stable all the time. However, the scalar modes in the small black hole (bottom left panel) exhibit anomalies with changes in temperature. When the temperature is lower than the critical point, the system is in a normal flow state, while above the critical temperature the system transits to a superfluid state. The right column is from the independent equation of the $b$ field. It shows an inflection point at the location of the phase transition with large black hole being our background geometry, but no inflection point appears in the case of a small black hole. Meanwhile, we explore the impact associated with chemical potential on the low-lying quasi-normal modes at fixed temperature, which is shown in Figure 9. Similarly, the dynamic behaviors of the low-lying quasi-normal modes undergoes a turning point at the phase transition location. Remarkably, the low-lying quasi-normal modes of $b$ field do not move until the phase transition occurs, which is reasonable according to Equation (25).
@@
Figure 8: The variation of the low-lying quasi-normal modes with the temperature at fixed chemical potential for large (top plots) and small (bottom plots) black holes, where angular quantum number is set to $l=1$. The temperature is varied from $T_{min}=\sqrt{3}/2\pi$ to 0.52 for large black hole, from $T_{min}=\sqrt{3}/2\pi$ to 0.743 for small black hole. As the temperature decreases, the two modes move in the direction indicated by the arrow. The left plots are the case of scalar channel and the right plots are from independent equation of $b$ field. The green dot represents phase transition from normal state to superfluid state for large black hole, the situation is exactly opposite for small black hole. The purple dot represents the state of the lowest temperature.
@@
Figure 9: The variation of the low-lying quasi-normal modes with chemical potential at fixed temperature for $l=1$ case under large (top plots) and small (bottom plots) black holes background. Chemical potential is varied from 3.5 to 5.0 for large black hole, from 2.2 to 3.5 for small black hole. Left plots are the case of scalar channel and the right plots are from independent equation of $b$ field. The green dot represents phase transition from normal state to superfluid state for both two types of black holes. Thus, we can identify that the modes from the $b$ field do not move until phase transition occurs.
@@
Furthermore, we investigate the impact of angular quantum numbers $l$ on collective modes. Figure 10 shows the dynamic behaviors of the collective modes with respect to the angular quantum number at fixed temperature and chemical potential for the normal-fluid state. Here, there exist three different channels of modes: the scalar field channel, the transverse gauge field channel, and the longitudinal gauge field channel threechannels . The three kinds of channels are decoupled when the system is in normal fluid state. In Figure 10, the three top panels are based on a large black hole background, and bottom results are from a small black hole background. The left column depict scalar field channel ($q_{1}$ and $q_{2}$ fields), middle column are from longitudinal gauge field channel (coupled $a$ and $c$ fields), and the right column are from transverse gauge field ($b$ field) channel. $l$ is varied from 0 to 6 for the scalar channel and from 1 to 6 for the gauge field channels. Since the fields of $b$ and $c$ are excitation modes of the angular components of the gauge fields, these excitation modes do not exist as $l=0$. Therefore, we calculate quasi-normal modes of $b$ field and $c$ field starting from $l=1$. Moreover, $a,c$ fields are coupled as $l\neq 0$ for normal-fluid state. There is only one zero mode for $a$ field, which reflects charge diffusion. Since $l=1$ on the sphere topology corresponds to the case of large wave number $k$ on the planar topology, the dynamic behaviors of the three channels for large black hole background in Figure 10 are similar to those in the planar case. However, we can identify that the dynamic behaviors of scalar modes in the backgrounds of the two types of black holes are completely different. Here, it also exhibits an unusual side for small black hole background. As such, compared to small black hole, the superfluid system with large black hole being background geometry in spherical topology is similar to the case that in planar topology. Figure 11 describes the superfluid case for different values of $l$. The left column is from scalar modes, the right column is from transverse gauge field. In this case, scalar field channel and longitudinal gauge field channel are coupled together, the transverse gauge field channel is always decoupled with other fields. Similarly, the dynamic behaviors of all modes under large black hole background in Figure 11 are consistent with those of the planar case, but there are anomalies in the case of small black holes.
@@
Figure 10: The variation of the low-lying quasi-normal modes with angular quantum number $l$ at fixed temperature and chemical potential within normal-fluid state for large (top panels) and small (bottom panels) black holes. The leftmost panels are about scalar field channel ($q_{1}$ field and $q_{2}$ field), middle panels are about longitudinal gauge field channel ($a$ field and $c$ field), the rightmost panels are about transverse gauge field ($b$ field). $l$ is varied from 0 to 6 for leftmost panels, from 1 to 6 for the rest.
@@
Figure 11: The variation of the low-lying quasi-normal modes with angular quantum number $l$ at fixed temperature and chemical potential within superfluid state for large (top plots) and small (bottom plots) black holes. The left plots are from scalar field modes, right plots are from transverse gauge field. $l$ is varied from 0 to 6 for top left panel, from 1 to 6 for the rest.